OpenStudy (lisa2345):

1. Jenny has a job that pays her $8 per hour plus tips, t. Jenny worked for 4 hours on Monday and made $65 in all. Which equation could be used to find t, the amount Jenny made in tips?

OpenStudy (afrodiddle):

what are the options?

OpenStudy (lisa2345):

I think the answer is 65=8(4)+t

OpenStudy (lisa2345):

am I right?

OpenStudy (afrodiddle):

no We are trying to find the tips

OpenStudy (afrodiddle):

t= 8(4) - 65

OpenStudy (lisa2345):

a. 65 = 4t + 8 b. 65= 8t / 4 c. 65 = 8t + 4 d. 65 = 8(4) + t

OpenStudy (lisa2345):

thats all the choices

OpenStudy (afrodiddle):

That's weird, then yes. you were right.

OpenStudy (afrodiddle):

@HWBUSTER00 That's kinda cheating, bruh.

OpenStudy (afrodiddle):

That's like Piratebay giving users access to pirated movies and games, still illegal.. still cheating.